# Reinforcement Learning with Promising Tokens for Large Language Models

### v1

@@ 1 Introduction
However, directly applying standard RL algorithms (e.g., PPO (Schulman et al., 2017), REINFORCE (Sutton et al., 1999)) to LLMs presents a unique challenge regarding the high dimensionality of the action space (Wen et al., 2024). Unlike traditional RL tasks with compact action spaces (e.g., Atari, Robot control (Mnih et al., 2013; Todorov et al., 2012)), the action space of an LLM corresponds to its full vocabulary, typically exceeding 50,000 tokens. In practice, previous work has attempted to alleviate this issue using inference strategies such as Top-$k$ (Hopkins et al., 2023) or Nucleus sampling (Holtzman et al., 2020), which prune the token space while maintaining text coherence during the rollout phase (Sheng et al., 2024). Yet, the policy optimization step still operates over the vocabulary size, creating a misalignment between efficient rollout and high-dimensional policy optimization. Consequently, the RL signal is distracted by the need to preserve basic syntactic coherence rather than focusing solely on the logical reasoning required for improvement.
To bridge this gap, we suggest that RL for LLM policies should operate within a more focused decision space, as illustrated in Fig. 1. We begin by investigating the distribution of valid reasoning steps and empirically verify that the tokens required for correct solutions are naturally concentrated within a small subspace of high-likelihood candidates. Building on this insight, we propose Reinforcement Learning with Promising Tokens (RLPT) method, which decouples logical decision-making from token generation. The core intuition of RLPT is to leverage the semantic priors of the pre-trained base model to identify a dynamic set of promising tokens, i.e., candidates with high likelihood that ensure syntactic correctness. Specifically, at each generation step, RLPT constructs a refined action space that restricts the RL process to optimize solely over these promising tokens via masking. By filtering out the long tail tokens, RLPT shifts the burden of the policy from predicting the next word to selecting the best move, allowing the agent to prioritize logical reasoning over the syntactic maintenance.
@@ 3.2 Analysis of Model-generated Solution
Table 2: Top-$K$ coverage rate (%) of successful trajectory tokens from model-generated solution.
@@ 6 Experiment
Method Math Domain Code Average
@@ Appendix B Omitted Proof
We analyze the variance of the gradient estimator $\hat{g}=\nabla_{\theta}\log\pi(a_{t}|s_{t})A_{t}$. We define the total variance of the gradient vector as the sum of element-wise variances (i.e., the trace of the covariance matrix), denoted as $\mathbb{V}(\hat{g})\triangleq\sum_{i\in\mathcal{V}}\text{Var}(\hat{g}_{i})$.
First, consider the gradient component for a single token $i$. Given the policy distribution $\pi(\cdot|s_{t})$, the indicator function $\mathbb{I}(a_{t}=i)$ is a Bernoulli variable with parameter $\pi_{i}=\pi(i|s_{t})$. Treating the advantage $A_{t}$ as a fixed scaling factor for the variance analysis of the policy noise, the variance of the $i$-th gradient component is:
@@
For standard optimization over the full vocabulary $\mathcal{V}$, the total variance is the sum over all tokens:
$\mathbb{V}(\hat{g}_{\text{full}})=\sum_{i\in\mathcal{V}}\pi_{i}(1-\pi_{i})A_{t}^{2}=\left(\sum_{i\in\mathcal{P}_{t}}\pi_{i}(1-\pi_{i})+\sum_{k\in\mathcal{T}_{t}}\pi_{k}(1-\pi_{k})\right)A_{t}^{2}.$ (8)
In RLPT, the policy is constrained to $\mathcal{P}_{t}$ and renormalized to $\tilde{\pi}$. The logits for tail tokens $k\in\mathcal{T}_{t}$ are masked, meaning their gradients are deterministically zero ($\text{Var}(\hat{g}_{\text{mask},k})=0$). The total variance for the masked policy is:
@@
Since the probability mass of the tail is negligible (as empirically verified), we verify the reduction under the approximation $\tilde{\pi}_{i}\approx\pi_{i}$ for $i\in\mathcal{P}_{t}$. The variance reduction is given by:
$\Delta\mathbb{V}=\mathbb{V}(\hat{g}_{\text{full}})-\mathbb{V}(\hat{g}_{\text{mask}})\approx\sum_{k\in\mathcal{T}_{t}}\pi_{k}(1-\pi_{k})A_{t}^{2}+\sum_{i\in\mathcal{P}_{t}}(\pi_{i}(1-\pi_{i})-\tilde{\pi}_{i}(1-\tilde{\pi}_{i}))A_{t}^{2}.$ (10)
Focusing on the dominant term from the removed tail tokens:
$\Delta\mathbb{V}\approx\sum_{k\in\mathcal{T}_{t}}\pi_{k}(1-\pi_{k})A_{t}^{2}.$ (11)
Due to the properties of the Softmax function, $\pi_{k}>0$ for all $k\in\mathcal{T}_{t}$. Assuming $A_{t}\neq 0$, the term $\pi_{k}(1-\pi_{k})A_{t}^{2}$ is strictly positive. This demonstrates that standard RL optimization over the full vocabulary inherently carries additional variance from the tail tokens. By constraining the action space, RLPT effectively eliminates this noise source. ∎
@@ C.1 Examples of the Datasets in Our Experiments
C.2 Prompts for LLMs
@@
  • Prompt for Code: Solve the following coding problem using the programming language python: {problems}

### v2

@@ 1 Introduction
However, directly applying standard RL algorithms (e.g., PPO (Schulman et al., 2017), REINFORCE (Sutton et al., 1999)) to LLMs presents a unique challenge regarding the high dimensionality of the action space (Wen et al., 2024). Unlike traditional RL tasks with compact action spaces (e.g., Atari, Robot control (Mnih et al., 2013; Todorov et al., 2012)), the action space of an LLM corresponds to its full vocabulary, typically exceeding 50,000 tokens. In practice, previous works have attempted to alleviate this issue using inference strategies such as Top-$k$ (Hopkins et al., 2023) or Nucleus sampling (Holtzman et al., 2020), which prune the token space while maintaining text coherence during the rollout phase (Sheng et al., 2024). Yet, the policy optimization step still operates over the vocabulary size, creating a off-policy mismatch between efficient rollout and high-dimensional policy optimization. Consequently, the RL signal is distracted by the need to maintain basic syntactic coherence rather than focusing solely on the logical reasoning required for improvement.
To bridge this gap, we suggest that RL for LLM policies should operate within a more focused decision space, as illustrated in Fig. 1. We begin by investigating the distribution of successful trajectories and empirically verify that the tokens required for correct solutions are naturally concentrated within a small subspace of high-likelihood candidates. Building on this insight, we propose Reinforcement Learning with Promising Tokens (RLPT) method, which decouples logical decision-making from token generation. The core intuition of RLPT is to enable the LLM policy to make decisions over the subspace of the token space, and focus more on strategic selection. To achieve this, RLPT leverages the semantic priors of the pre-trained base model to identify a dynamic set of promising tokens, i.e., candidates with high likelihood that ensure syntactic correctness. At each generation step, RLPT constructs a refined action space that restricts the RL process to optimize solely over these promising tokens via masking. By filtering out the long tail tokens, RLPT shifts the burden of the policy from predicting the next word to selecting the best move, allowing the agent to prioritize logical reasoning over the syntactic maintenance.
@@ 3.2 Analysis of Model-generated Solution
Table 2: Top-$K$ coverage rate (%) of successful trajectory tokens from model-generated solutions.
@@ 6 Experiment
Method Math Telecom Code Average
@@ Appendix B Omitted Proof
First, we calculate the variance of the gradient component for a single token $i$. Since $a_{t}\sim\pi(\cdot|s_{t})$, the term $\mathbb{I}(a_{t}=i)$ is a Bernoulli variable with parameter $\pi_{i}=\pi(i|s_{t})$. The variance of $\hat{g}_{i}$ is:
@@
For the standard optimization over the full vocabulary $\mathcal{V}$, the total variance sums over all tokens:
$\mathbb{V}(\hat{g}_{\text{full}})=\sum_{i\in\mathcal{V}}\pi_{i}(1-\pi_{i})A_{t}^{2}=\sum_{i\in\mathcal{P}_{t}}\pi_{i}(1-\pi_{i})A_{t}^{2}+\sum_{k\in\mathcal{T}_{t}}\pi_{k}(1-\pi_{k})A_{t}^{2}.$ (8)
In RLPT, the policy is renormalized to $\tilde{\pi}$ over $\mathcal{P}_{t}$, and the logits for tail tokens $k\in\mathcal{T}_{t}$ are masked (i.e., gradients are deterministically zero). Thus, $\text{Var}(\hat{g}_{\text{mask},k})=0$ for all $k\in\mathcal{T}_{t}$. The total variance becomes:
@@
Since the tail probability mass is negligible (as empirically verified), we have $\tilde{\pi}_{i}\approx\pi_{i}$ for $i\in\mathcal{P}_{t}$. The reduction in variance is dominated by the removal of the tail sum:
$\Delta\mathbb{V}=\mathbb{V}(\hat{g}_{\text{full}})-\mathbb{V}(\hat{g}_{\text{mask}})\approx\sum_{k\in\mathcal{T}_{t}}\pi_{k}(1-\pi_{k})A_{t}^{2}.$ (10)
Since $\pi_{k}>0$ for tokens in the tail (due to the Softmax properties) and $A_{t}^{2}\geq 0$, the term $\sum_{k\in\mathcal{T}_{t}}\pi_{k}(1-\pi_{k})A_{t}^{2}$ represents a strictly positive quantity. This explicitly shows that standard RL optimization suffers from noise injection caused by the gradients of thousands of irrelevant tail tokens, which RLPT effectively eliminates. ∎
@@ C.1 Examples of the Datasets in Our Experiments
C.2 Prompts Used in Our Experiments
Tab. 7 presents the prompts used in our exeriments.
Table 7: Prompts for Experiments.
Task Name Prompt
@@ C.2 Prompts Used in Our Experiments
Code Solve the following coding problem using the programming language python: {problems}